OpenStudy (anonymous):

Is it possible to solve this problem? 1*g(0) + 1*g(1) + 3*g(-1) + 3*g(2) = 0 I know the function 0(x) would work, but, is there a way to find another polynomial with a degree of at most 3 that satisfies this equation?

OpenStudy (anonymous):

I tried to make is so that it was 1 -1 3 -1 and I got a function g(x)=-2x+1. Is that a valid solution do you think? Is there a mathematical way to actually do this? I just guessed and checked, and got lucky! aha

OpenStudy (mertsj):

I don't think I understand the problem.

OpenStudy (anonymous):

I have this weird inner product space defined as f*g = f(0)g(0) + f(1)g(1) + f(-1)g(-1) + f(2)g(2) and I am given v=x^2 -x +1 I want to find a polynomial w, such that v*w = 0 (the * as defined in f*g) so, I found f(0)=1 f(1)=1 f(-1)=3 f(2)=3 so, now my f*g becomes 1g(0) + 1g(1) +3g(-1) + 3g(2)= 0 I want to find this function g so that this holds true ... (it turns out that g(x)=-2x +1 fits this definition. But this was just a lucky guess)

OpenStudy (anonymous):

so, I am saying that w= -2x+1. If this is confusing you, you're in the same boat. I'm pretty confused myself about it. I hope I am doing it correctly!